OpenStudy (anonymous):

There are two similar cones. Cone a has an altitude of 5cm and cone b has an altitude of 13cm. If cone a weighs 20lbs, how much does cone b weigh? Enter in decimal form, no rounding.

OpenStudy (anonymous):

Is the volume of a cube \[\pi \times Altitude \times R ^{3}\]

OpenStudy (anonymous):

*sorry not cube, CONE

OpenStudy (anonymous):

oh, i forgot the 1/3

OpenStudy (anonymous):

351.52 lbs

OpenStudy (yamyam70):

Don't just give out the answers, show your work.

OpenStudy (anonymous):

no problem

OpenStudy (anonymous):

Weight = mass * g_constant mass = density * volume this is really a problem about finding the volume of a cone I derived the volume as \[V = ( \pi r ^{2}h) \div 3\]

OpenStudy (anonymous):

Since cone b is 2.6 times taller than cone a, the radius of the base of cone b is 2.6 times the base radius of cone a. using the volume formula \[R_b = 2.6 \times \sqrt{(20 lbs \times 3) \div (5cm \times \pi \times \delta \times g) }\] where R_b is the radius at the base of cone b, delta is density and g is the gravitational constant. Plug R_b into the volume equation and multiply it by delta and g to get your answer. 351.52 lbs

OpenStudy (yamyam70):

:) @amber1227 I hope you get that :) , welcome to open study! :)